ABC Pvt. Ltd. is considering two mutually exclusive capital investments. The project’s expected net cash flows are as follows:
Expected Cash Flows
Year Project A Project B
0 -375 -575
1 -300 190
2 -200 190
3 -100 190
4 600 190
5 600 190
6 926 190
7 -200 0
If you were told that each project’s cost of capital was 12%, which project should be selected using the NPV criteria? What is each project’s IRR? What is the regular payback period for these two projects? What is the profitability index for each project if the cost of capital is 12%?
1
Expert's answer
2020-12-10T10:06:19-0500


Net present value of project (B)

NPV = -initial cost + ACF(PVIF12%,6)

NPV = -575 + 190(PVIF12%,6)

NPV "= - 575 +190 \times 4.1114"

NPV = $206.17

Based on NPV Creterian Project (A) should be selected, because of its higher NPV

Calculation of IRR of both the projects, using Excel function of IRR would be:

Project A = 18.64%

Project B = 23.92 %

Regular payback periods:

Project A = 4 years + "\frac{375}{600}" = 4.625 years

Project B = "\frac{575}{190}" = 3.03 years

Profitability index (PI):

PI = PV cash inflow/initial costs

PI (A) = "\frac{601.90}{375} = 1.6051"

PI (B) = "\frac{781.17}{575} = 1.3586"